If V=6 and E=9, find F using Euler’s formula. What is euler’s formulae?

Grade:6

1 Answers

T.B.S.A T.B.S.A
5 years ago
V-E+F=2 Is Euler’s formula  So 6-9+F=2 So F=5 So -3+F=2 So F=3+2 So it becomes F=5
So if there is 6vertices ,9edges then there is 5faces